# MeshUp: Multi-Target Mesh Deformation via Blended Score Distillation

### v2

@@ Abstract
We propose MeshUp, a technique that deforms a 3D mesh towards multiple target concepts, and intuitively controls the region where each concept is expressed. Conveniently, the concepts can be defined as either text queries, e.g., “a dog” and “a turtle,” or inspirational images, and the local regions can be selected as any number of vertices on the mesh. We can effectively control the influence of the concepts and mix them together using a novel score distillation approach, referred to as the Blended Score Distillation (BSD). BSD operates on each attention layer of the denoising U-Net of a diffusion model as it extracts and injects the per-objective activations into a unified denoising pipeline from which the deformation gradients are calculated. To localize the expression of these activations, we create a probabilistic Region of Interest (ROI) map on the surface of the mesh, and turn it into 3D-consistent masks that we use to control the expression of these activations. We demonstrate the effectiveness of BSD empirically and show that it can deform various meshes towards multiple objectives. Our project page is at this URL
@@ 2 Related Work
Several data-driven techniques have been used to learn deformations [1, 20, 66, 37, 19, 58, 15]. A recent class of works leverages text prompts as user inputs for driving a deformation towards an arbitrary textual prompt [38, 28, 18]. These methods use various deformation representations, and we opt to leverage Jacobians since they produce smooth and large-scale global deformations.
@@ 3 Method
Using this SDS gradient, one can deform a mesh to a single target prompt. A detailed derivation could be found in the supplementary materials.
@@
For clarity, let j denote the index for the $j^{th}$ target-branch, each associated with a target ”concept.” The $j^{th}$ branch would take as input its associated target text, $y_{j}$, and a weight $w_{j}$ that controls the degree to which $y_{j}$ should be expressed. The key observation is that the activation matrices ($\phi^{j}$) we get at the end of each attention layer represent the“weighted feature space” of each concept, defined over the space of the patch of the input renderings. To blend the features over this space, we perform a weighted interpolation of the activations across the patch dimension, and inject them into the corresponding patch location in the blending branch. Formally, we inject the activation for a single concept as follows:
$\phi^{blend}\leftarrow w_{j}\dot{\phi}^{j}+(1-w_{j})\dot{\phi}^{blend}$ (5)
@@ 4 Experiments
In this section, we first show multi-target deformation results driven by text or image targets. Additionally, we demonstrate deformations with local control and mesh targets. Finally, we also describe how our method can be used as a regularization term that controls the strength of a deformation. We provide comparisons with various baselines, show an experiment that uses our method to perform keypoint interpolation between concepts, and show a qualitative user study of our method in the supplementary material.
@@ 5 Conclusion and Limitations
In this paper, we propose a versatile framework for mesh deformation that supports creative workflows, enabling deformation via text or image-based concepts, mixing these concepts using various weights, and localizing their expressions.
@@
Technical Details In this section, we will go over some of the technical details of our model.
@@
First, we note that we use the IF model (XL size) by Deepfloyd for our denoiser. While the IF model consists of three components–a Text-to-Image Diffusion model (Unet) and two super-resolution models–we only use its Text-to-Image Denoiser. As opposed to the widely used Stable-Diffusion that operates in the latent embedding space, IF does so in the rgb space, and reducing both memory usage and training time, without compromising quality. In addition to the denoiser, the t5 text encoder on which IF is trained has a much more expressive embedding space than that of CLIP. Since we aim to employ Textual Inversion to learn a set of images through this embedding space, the t5 encoder comes to our great advantage.
@@
We ran our results using a single A40 gpu, and ran 2400 iterations with a batch size (number of renderings per iteration) of 16. It takes approximately 1.5 hours to optimize for a single target. The time increases by a factor of approximately 1.5 for each target that we add for blending. However, we have observed that the deformation results get reasonably similar to the final results after 500-600 iterations, which takes about 20 minutes to run for a single target. Running the localized deformation method also increases the run-time by an approximate factor of 1.5, not including the optional LoRA-finetuning stage, which takes about an extra 10-15 minutes. The optimization time also depends on the number of vertices/faces of the source mesh, and we can therefore only provide a rough approximation of the run-time.
@@
We also note that we use the nvdiffrast model to rasterize the mesh into 2D images from multiple viewpoints. Since our main objective is to optimize the geometry of the source mesh (and not its texture), we simply paint the mesh with a uniformly grey texture, and rasterize it to get the input renderings. We found that the grey texture works best for our purpose, but we leave the question of improving the geometry with a more sophisticated texturing scheme up for future work. We first rasterize the mesh in 512x512 resolution and downsample it to 64x64 using bilinear interpolation before feeding it to the diffusion model.
@@
Specific Overview of SDS We will go over the SDS perspective of diffusion in detail, and how we apply this perspective for our purpose of deforming the Jacobians. We first limit our scope to 2D image-to-image generation, where, given an input image $\mathbf{z}$, the objective is to optimize some parameter $\phi$ in the 2D space. Given a text condition $\mathbf{t}$, we define the loss function to be the squared L2 norm between the noise predicted by the denoising model $\epsilon_{\phi}$ and the sampled noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$.
@@
Note $t$ is the timestep randomly sampled from a uniform distribution, $t\sim\mathcal{U}(0,1)$, $\epsilon$ is a random noise sampled from a standard normal distribution, and $\mathbf{z}_{\mathbf{t}}$ is the input image noised for timestep $t$ using the re-parameterization trick whereby we conveniently get the noise at timestep $\mathbf{t}$ using the equation, $\mathbf{z}_{\mathbf{t}}=\sqrt{\alpha_{t}}\mathbf{z}+\sqrt{1-\alpha_{t}}\epsilon$. $w(t)$ is a weighting function for which we will not go into details. In simple terms, we aim to optimize some parameters so that the (frozen) model can precisely predict the noise sampled from timestep $t$. When this loss is minimized, the parameters are optimized to represent an object (or an image rendered from the object) that is as close as possible to that ”guided” by the denoiser.
To optimize for some parameter $\phi$, which, in most image generation applications, will be the image pixel value, we guide $\phi$ toward the highest probability region predicted by the denoiser.
@@
The gradient of $\mathcal{L}_{\text{Diff}}$ with respect to $\phi$, which we denote $\nabla_{\theta}\mathcal{L}_{\text{Diff}}$, can be derived by,
@@
$\nabla_{\theta}\mathcal{L}_{\text{Diff}}=\left(\epsilon_{\phi}\left(\mathbf{z}_{\mathbf{t}},y,t\right)-\epsilon\right)\frac{\partial\epsilon_{\phi}(\mathbf{z},y,t)}{\partial\mathbf{z}_{\mathbf{t}}}\frac{\partial\mathbf{z}_{\mathbf{t}}}{\partial\theta}.$ (12)
@@
It is known from [44] that instead of having to backpropagate through the denoiser we can approximate an effective gradient for $\phi$, denoted as $\nabla_{\theta}\mathcal{L}_{\mathrm{SDS}}$, simply by omitting the gradient with respect to the denoiser, $\frac{\partial\epsilon_{\phi}(\mathbf{z},y,t)}{\partial\mathbf{z}_{\mathbf{t}}}$, giving us the equation,
@@
$\nabla_{\theta}\mathcal{L}_{\mathrm{SDS}}(\mathbf{z},y,\epsilon,t)=(\epsilon_{\phi}\left(\mathbf{z}_{\mathbf{t}},y,t\right)-\epsilon)\frac{\partial\mathbf{z}_{\mathbf{t}}}{\partial\theta}$ (13)
@@
In addition to SDS, we also use classifier-free guidance with an extremely high weight of 100 to help with our 3D objective, following the observation from [44].
Using SDS to guide Mesh Deformation In order to optimize for the Jacobians $\mathbf{J_{i}}$ of mesh faces instead of some arbitrary parameter $\phi$, we can simply multiply $\nabla_{\mathbf{J_{i}}}\mathcal{L}_{\mathrm{SDS}}$ to any additional gradients involved in calculating the image renderings from the optimized per-face Jacobians. We can change our equation into,
@@
$\nabla_{\mathbf{J_{i}}}\mathcal{L}_{\mathrm{SDS}}=(\epsilon_{\phi}\left(\mathbf{z}_{\mathbf{t}},y,t\right)-\epsilon)\frac{\partial\mathbf{G}(\mathbf{J_{i}})}{\partial\mathbf{J_{i}}}.$
@@
where we abuse notation and abstractly denote $\mathbf{G}(\mathbf{J_{i}})$ as the span of the entire computational pipeline, ranging from the optimized 3x3 Jacobians to the renderings of the deformed mesh (this includes the operations of projecting the 3x3 Jacobians to the 3x2 space, running the poison solve to calculate the deformation map $\phi$, deforming the mesh, rendering this deformed mesh onto the image space). In practice, we use Pytorch’s autograd library to automatically handle the differentiation of $\mathbf{G}(\mathbf{J_{i}})$.
@@
Localized Deformation We will now detail the implementation of our localized deformation experiment.
@@
First, we note that each attention layer of the Unet denoiser takes as input the text encoding and the hidden states, or activations, given by the previous attention layer. We then project the hidden states each onto the embedding space of Key, Query, and Value. These projected matrices are then concatenated with the similarly projected text encoding vector, giving us the Key, Query, and Value matrices. [7, 59] Specifically, the attention map is defined as
@@
Note that the attention matrix $M$ contains both the self- and cross-attention map, where the cross-attention maps the correlation between the text prompt embeddings and various ”patches” (or regions) within the image and the self-attention map does so between the ”patches” themselves. Using this observation, we utilize the self-attention maps to identify the probable regions that have high correspondence with the patches to which the control vertices are mapped. Once we extract the 3D-consistent ROI mask $R^{\prime}_{m}$ using our localized control method, we apply this mask to the cross-attention map of each target, thereby masking out the regions within the image where the target concepts (represented by the text prompt embeddings) can be expressed.
We will now delineate some of the implementation details of this localized deformation method. First, as briefly mentioned in the method section, we fine-tune the denoising Unet of our DeepFloyd IF model to make it precisely predict the sampled noise from the renderings of the input mesh. In other words, we train the model such that it can precisely reconstruct the renderings during the inference time. The objective of our fine-tuning is similar to that proposed in  [47], but we use multi-viewpoint mesh renderings as inputs and do not use the Class Prior Preservation Loss. This fine-tuning has a similar effect to inverting the renderings of the mesh, encouraging the model to be capable of reconstructing the mesh renderings.
Moreover, since fine-tuning the entire U-net as implemented in the original Dreambooth paper is prohibitively expensive, we instead finetune the LoRA weights  [26] with the same training objective as described above. We run 150 fine-tuning iterations on a batch of 16 images, and the entire operation takes approximately 15 mins.
@@
Additionally, we also use the fine-tuned weights to stabilize the localized deformation method. Notably, when the weight $w_{j}$ is particularly high (over 0.8) for a certain target, the activation $\phi^{blend}=w_{i}\phi^{i}+w_{j}\phi^{j}$ might become extremely unstable as $w_{i}$ could be too small so that the activation $\phi^{blend}$ receives extremely low amount of signal, or weights, from the region masked-out from the $j^{th}$ localization targets. In order to compensate for this instability, whenever the $(1-{max}(w_{j},w_{i}))<0.2$ (when one target has a dominently high weight), we create an inverse boolean mask of $R^{\prime}_{m}$, ${}R^{\prime}_{m}$ and apply this mask to the cross-attention from the LoRA-finetuned text-token. We then weight this cross attention by $w_{lora}=0.2$, and add it to the cross-attention mask of the $j^{t}h$ target (the target within $w_{j}>0.8)$. This operation ensures that all the regions of the mesh receive at least a minimal amount of attention from all regions required to keep the activation $\phi^{j}$ stable throughout the deformation process.
The entire localization process takes about 2 hours to complete (for a 2-way blending of concepts) but could be reduced by a factor of approximately 1.3 by extracting the self-attention maps $R_{m}$ from the original diffusion processes conditioned on the target concepts without fine-tuning the LoRA weights. The quality of localization, however, could be compromised.
@@
Modified Classifier Free Guidance for Self-Blending In this section, we will delineate the modified classifier free guidance (CFG) method that we use for our self-blending experiments. In the self-blending experiment, we reduce our pipeline to take a single concept as input and control the weights $w_{j}$ to regularize the expression of the concept. We therefore reduce the branches of the multi-target pipeline into one deformation and a target branch and run two parallel pipelines instead of many.
One caveat to this method, however, is that when the target branches are reduced to one, there is only one text/image target that can replace the target encodings of the deformation branch, which always leaves the deformation branch partially conditioned on the null text prompt ””, introducing some undesired bias for $\nabla_{theta}L_{SDS}$.
To counter this problem, we slightly modify our equation for classifier-free guidance as follows. The original Classifier Free Guidance, a method introduced by [23] to alleviate the bias and complexity entangled with the classifier inputs of diffusion models, is formulated as follows,
@@
where $\hat{\epsilon}_{text}$ and $\hat{\epsilon}_{null}$ are the predicted noise conditioned on the target text prompt and null text prompt, respectively. $\alpha$ is the Guidance Scale, a parameter that defines the strength of the classifier-free guidance:
@@
where we replace $\hat{\epsilon}_{text}$ with $\epsilon$, the sampled noise.
@@
Such modification to the CFG score allows our model to achieve a more unbiased control of the deformation strength, as opposed to the original CFG. More intuitively, we want there to be no deformation when the weight is set to 0. The modified CFG ensures that the gradient $\nabla_{\theta}L_{\text{Diff}}=0$ when we set the deformation strength to 0, since if we do not inject any attention from the target branch, the predicted noise becomes:
@@
Injecting attention to this modified CFG score ensures a more stable control of the deformation strength by getting rid of the bias that $\epsilon_{null}$ introduces on the score. The modified CFG score thereby provides guidance that aligns more closely with the intuition of interpolating between the ”identity deformation” and ”deformation conditioned on the target prompt.”
We additionally note that we can replace $\hat{\epsilon}_{text}$ with $\epsilon$, with minimal sacrifice in quality because we use a significantly higher guidance scale $\alpha$ of 100, following the findings of  [44]. Such a high guidance scale makes $\hat{\epsilon_{text}}$ relatively insignificant compared to the Classifier Free Guidance term, ensuring a minimal sacrifice in quality.
Appendix A Additional Experiments
Continuous Concept Space. Finally, users might want to explore a continuous space of generated concepts. While we could run our pipeline multiple times with different relative weights between targets, this could easily become prohibitively expensive if users want a smooth continuity (e.g., generating morphing animations) and have a large combination of targets. To address this challenge, we observe that our mesh-based representation provides a dense correspondence map between the source and the deformed shapes. Thus, once the user generates a few sample key frame shapes, they can be smoothly interpolated continuously, offering a powerful space for concept exploration.

### v3

@@ Abstract
We propose MeshUp, a technique that deforms a 3D mesh towards multiple target concepts, and intuitively controls the region of their expression. Conveniently, the concepts can be defined as either text queries, e.g., “a dog” and “a turtle,” or inspirational images, and the local regions can be selected as any number of vertices on the mesh. We can effectively control the influence of the concepts and mix them together using a novel score distillation approach, referred to as the Blended Score Distillation (BSD). BSD operates on each attention layer of the denoising U-Net of a diffusion model as it extracts and injects the per-objective activations into a unified denoising pipeline from which the deformation gradients are calculated. To localize the expression of these activations, we create a probabilistic Region of Interest (ROI) map on the surface of the mesh, and turn it into 3D-consistent masks that we use to control the expression of these activations. We demonstrate the effectiveness of BSD empirically and show that it can deform various meshes towards multiple objectives. Our project page is at https://threedle.github.io/meshup/.
@@ 2 Related Work
Several data-driven techniques have been used to learn deformations [1, 20, 68, 38, 19, 59, 15, 33]. A recent class of works leverages text prompts as user inputs for driving a deformation towards an arbitrary textual prompt [39, 28, 18, 66]. These methods use various deformation representations, and we opt to leverage Jacobians since they produce smooth and large-scale global deformations.
@@ 3 Method
Using this SDS gradient, one can deform a mesh to a single target prompt. A detailed discussion can be found in the supplementary materials.
@@
For clarity, let j denote the index for the $j^{th}$ target-branch, each associated with a target “concept.” The $j^{th}$ branch would take as input its associated target text, $y_{j}$, and a weight $w_{j}$ that controls the degree to which $y_{j}$ should be expressed. The key observation is that the activation matrices ($\phi^{j}$) we get at the end of each attention layer represent the“weighted feature space” of each concept, defined over the space of the patch of the input renderings. To blend the features over this space, we perform a weighted interpolation of the activations across the patch dimension, and inject them into the corresponding patch location in the blending branch. Formally, we inject the activation for a single concept as follows:
$\phi^{blend}\leftarrow w_{j}\phi^{j}+(1-w_{j})\phi^{blend}$ (5)
@@ 4 Experiments
In this section, we first show multi-target deformation results driven by text or image targets. Additionally, we demonstrate deformations with local control and mesh targets. Finally, we also describe how our method can be used as a regularization term that controls the strength of a deformation. We provide comparisons with various baselines, show an experiment that uses our method to perform keypoint interpolation between concepts, and show a qualitative user study of our method in the supplementary material. We also provide a comprehensive collection of various deformation results in Figure 15.
@@ 5 Conclusion and Limitations
In this paper, we propose a versatile framework for mesh deformation that supports deformation via text or image-based concepts, mixing these concepts using various weights, and localizing their expressions.
@@
The following sections will provide more details of our method, present additional experiments, and evaluate our method using quantitive/qualitative metrics. Section A will delineate the technical and mathematical details of our method, including model configurations, runtime, Score Distillation Sampling (SDS), localized deformation, and the modified classifier free guidance score used in our self-blending experiments. Section B will present the experiments, and Section C will provide evaluations.
Appendix A Technical Details
@@
First, we briefly justify our choice of DeepFloyd-IF XL model as our denoiser. For our purposes, we noticed that IF performs significantly better than the widely used latent Diffusion models, most likely as it operates directly in the rgb space. In addition to the denoiser, the t5 text encoder on which IF is trained has a much more expressive embedding space than that of CLIP. Since we aim to employ Textual Inversion to learn a set of images through this embedding space, the t5 encoder comes to our great advantage. We leave more extensive comparison across various diffusion models to future work.
@@
For runtime, we ran all of our results using a single A40 gpu, and ran 2400 iterations with a batch size (number of renderings per iteration) of 16. It takes approximately 1.5 hours to optimize for a single target. The time increases by a factor of approximately 1.5 for each target that we add for blending. However, we have observed that the deformation converges reasonably close to the final results after 600-800 iterations, which takes about 25-30 minutes to run for a single target. Running the localized deformation method also increases the run-time approximately by an factor of 1.5, not including the optional LoRA-finetuning stage, which takes an extra 10-15 minutes to converge. The optimization time also depends heavily on the number of faces of the source mesh, and we can therefore only provide a rough approximation of the run-time. Most of the meshes that we used in this paper had 10,000-30,000 faces.
@@
For rendering, we use nvdiffrast to rasterize the mesh from multiple viewpoints. Since our main objective is to optimize the geometry of the source mesh (and not its texture), we simply paint the mesh with a uniformly grey (0.5, 0.5, 0.5) texture, and rasterize it to get the renderings. We found the grey texture works reasonably well for our purpose, but we leave the investigation of more sophisticated texturing schemes up for future work. We also rasterize the mesh in 512x512 resolution and downsample it to 64x64 using bilinear interpolation before inputing it to the diffusion model.
@@
Specific Overview of SDS We will go over the SDS perspective of diffusion in detail, and how we apply this perspective for our purpose of deforming the Jacobians. We first limit our scope to 2D image-to-image generation, where, given an input image $\mathbf{z}$, the objective is to optimize some parameter $\theta$ in the 2D space. Given a text condition $\mathbf{t}$, we define the loss function to be the squared L2 norm between the noise predicted by the denoising model $\epsilon_{\phi}$ and the sampled noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$:
@@
where $t$ is the randomly sampled timestep from a uniform distribution, $t\sim\mathcal{U}(0,1)$, $\epsilon$ is a random noise sampled from a standard normal distribution, and $\mathbf{z}_{\mathbf{t}}$ is the input image noised according to timestep $t$ using the re-parameterization trick, $\mathbf{z}_{\mathbf{t}}=\sqrt{\alpha_{t}}\mathbf{z}+\sqrt{1-\alpha_{t}}\epsilon$. $w(t)$ is a weighting function for which we will not go into details. In simple terms, we aim to optimize some parameters so that the (frozen) model can precisely predict the noise sampled from timestep $t$. When this loss is minimized, the parameters ($\theta$) are optimized to represent an object that is as close as possible to what is predicted by the denoiser to be of highest probablity. A deeper analysis of diffusion models and Score Distillation Sampling can be found in [24, 61, 45]
@@
The gradient of $\mathcal{L}_{\text{Diff}}$ with respect to $\theta$, which we denote $\nabla_{\theta}\mathcal{L}_{\text{Diff}}$, can be derived by,
@@
$\nabla_{\theta}\mathcal{L}_{\text{Diff}}=\left(\epsilon_{\phi}\left(\mathbf{z}_{\mathbf{t}},y,t\right)-\epsilon\right)\frac{\partial\epsilon_{\phi}(\mathbf{z},y,t)}{\partial\mathbf{z}_{\mathbf{t}}}\frac{\partial\mathbf{z}_{\mathbf{t}}(\theta)}{\partial\theta}.$ (12)
@@
It is known from [45] that instead of having to backpropagate through the denoiser we can approximate an effective gradient for $\theta$, denoted as $\nabla_{\theta}\mathcal{L}_{\mathrm{SDS}}$, simply by omitting the gradient with respect to the denoiser, $\frac{\partial\epsilon_{\phi}(\mathbf{z},y,t)}{\partial\mathbf{z}_{\mathbf{t}}}$, giving us,
@@
$\nabla_{\theta}\mathcal{L}_{\mathrm{SDS}}(\mathbf{z},y,\epsilon,t)=(\epsilon_{\phi}\left(\mathbf{z}_{\mathbf{t}},y,t\right)-\epsilon)\frac{\partial\mathbf{z}_{\mathbf{t}}(\theta)}{\partial\theta}$ (13)
@@
In addition to SDS, we also use classifier-free guidance with an extremely high classifier-free guidance weight of 100 to help with our 3D objective [45].
Using SDS to guide Mesh Deformation In order to optimize for the Jacobians $\mathit{J_{i}}$ of mesh faces instead of some arbitrary parameter $\theta$, we can simply replace $\theta$ in (13) with $\mathit{J_{i}}$,
@@
$\nabla_{\mathbf{J_{i}}}\mathcal{L}_{\mathrm{SDS}}=(\epsilon_{\phi}\left(\mathbf{z}_{\mathbf{t}},y,t\right)-\epsilon)\frac{\partial\mathbf{z}_{\mathbf{t}}(\mathbf{J_{i}})}{\partial\mathbf{J_{i}}}.$
@@
We denote $\mathbf{z}_{\mathbf{t}}$ to be dependent on $\theta$ since we denote it to encompass the span of the entire computational pipeline, from the optimized 3x3 Jacobians to the renderings of the deformed mesh (this includes the operations of projecting the 3x3 Jacobians to the 3x2 space, running the poison solve to calculate the deformation map $\phi$, deforming the mesh, rendering this deformed mesh onto the image space). In practice, we use Pytorch’s autograd library to automatically handle the differentiation of $\mathbf{z}_{\mathbf{t}}(\mathbf{J_{i}})$.
@@
Localized Deformation We will now detail the implementation of our localized deformation experiment. This explanation is specific to the DeepFloyd IF model that we use
@@
First, we note that the UNet denoiser consists of multiple attention layers, and each of these layer takes as input the text encoding and the hidden states, output by the previous layer. We then project the hidden states each as Key, Query, and Value matrices using learned mlp layers. These matrices are then concatenated with matrices similarly projected from the text enbeddings, giving us the matrices, $Q$, $K$, $V$. [7, 60] Specifically, the attention map is defined as
@@
The attention matrix $M$ consists of both the self- and cross-attention map. The cross-attention map encapsulates the correlation between the text embeddings and the various “patches” of the image, while the self-attention map finds the correlation across these patches. Using this observation, we first utilize the self-attention maps to identify the regions on the image that have high correspondence with the control vertices, and convert this map into a 2D Region of Interest (ROI) mask, $R_{m}$. We then extract the 3D-consistent ROI mask $R^{\prime}_{m}$ using our rasterization function, project it back onto 2D, and use it to mask out the cross-attention map of each target. Effectively, we are distilling the regions within the image where target concepts should be expressed.
We will now delineate some of the implementation details of the localized deformation method. First, as briefly mentioned in the method section, we aim to replicate the effect of inverting the renderings by finetune the denoising UNet of our DeepFloyd IF model by the objective proposed in  [48]. While dreambooth finetunes the denoiser so that it can reproduce the input image, we do so for multi-viewpoint mesh renderings, encouraging the model to be capable of reconstructing the mesh renderings given a predefined token. Moreover, since fine-tuning the entire UNet as implemented in the original Dreambooth paper is prohibitively expensive, we instead finetune the LoRA weights  [26] with the same training objective as described above. We run 150 fine-tuning iterations on a batch of 16 images, and the entire operation takes approximately 15 mins. While this finetuning process betters the quality of the localization map, we also observe that the map is reasonable enough even without finetuning. Users who desire faster generation time over better localization could skip this process with minimal compromise in quality.
@@
Additionally, we use the finetuned weights to stabilize the localized deformation method. Notably, when a particular weight $w_{i}$ is extremely low (over 0.2) for a certain target, the blended activation $\phi^{blend}=w_{i}\phi^{i}+w_{j}\phi^{j}$ (where $i$ and $j$ each represent different target concepts) might become extremely unstable as the localized region corresponding to the $i^{th}$ concept could only receive minimal activation signals. In order to compensate for this instability, whenever the $(1-{max}(w_{j},w_{i}))<0.2$ (when one target has extremely small weight), we create an inverse boolean mask, $\sim R^{\prime}_{m}$ and apply this mask to the cross attention map of the finetuned UNet. We then weight this masked cross attention by $w_{lora}=0.8$, and simply add it to the cross-attention mask of the $i^{t}h$ target. This operation ensures that all the regions of the mesh receive at least a minimal amount of attention required to keep the activation $\phi^{j}$ stable throughout the deformation process.
With the localization process, the entire deformation process takes about 2 hours to complete (for a 2-way blending of concepts).
@@
Modified Classifier Free Guidance for Self-Blending In this section, we will discuss how we modified the Classifier Free Guidance (CFG) for our self-blending experiments (Figure 11). In the self-blending experiment, we utilize our blending pipeline to control the expression of a single concept. We thus reduce the branches to two (one target branch and one blending branch), and extract the activation as $\phi^{blend}=w\phi+(1-w)\phi^{null}$, where $\phi^{null}$ is the activation from the blending branch (activation from the null text promt, “”).
One caveat to this method, however, is that $\phi^{null}$ could introduce undesired bias for $\nabla_{theta}L_{SDS}$, especially when its corresponding weight, $1-w$, is high.
To tackle this problem, we slightly modify our equation for CFG The original CFG, a method introduced by [23], is formulated as follows:
@@
where $\hat{\epsilon}_{text}$ and $\hat{\epsilon}_{null}$ are the predicted noise conditioned on the target text prompt and null text prompt, respectively. $\alpha$ is the Guidance Scale, a parameter that controls the strength of CFG. We modify our equation into
@@
where we simply replace $\hat{\epsilon}_{text}$ with $\epsilon$, the sampled noise.
@@
Such modification allows our model to achieve a more stable, unbiased control of the deformation strength, as opposed to the original CFG, as shown in Figure 18. Intuitively, we want the result to identity (no deformation) when the weight is set to 0, which is precisely what the modified equation achieves. As follows, the modified CFG ensures that the gradient $\nabla_{\theta}L_{\text{Diff}}=0$ when the deformation strength is 0, where our equation gives,
@@
Injecting attention to this modified CFG score ensures a more stablized control of the deformation strength by getting rid of the bias introduced by $\epsilon_{null}$. The modified CFG score thereby provides guidance that aligns more closely with the intuition of interpolating between the “identity deformation” and “deformation conditioned on the target prompt.”
We additionally note we can replace $\hat{\epsilon}_{text}$ with $\epsilon$, with minimal compromise in quality because we use a significantly high guidance scale $\alpha$ of 100, following the findings of  [45]. Such a high guidance scale makes $\hat{\epsilon_{text}}$ relatively insignificant compared to the Classifier Free Guidance term, ensuring minimal change in quality.
Appendix B Additional Experiments
Continuous Concept Space. Here, we explore a scenario when our users might want to explore a continuous space of generated concepts. While we could run our pipeline multiple times with different relative weights between targets, this could easily become prohibitively expensive if users want a smooth continuity (e.g., generating morphing animations). To address this challenge, we observe that our mesh-based representation provides a dense correspondence map between the source and the deformed shapes. Thus, the user could generate a few sample key frame shapes using our method and smoothly interpolated between them. We show this experiment in Figure 16